# MMGPL: Multimodal Medical Data Analysis with Graph Prompt Learning

### v1

@@ III-D2 Graph embedding
where $\tilde{\mathbf{T}}$ is the weighted embeddings of the tokens. As a result, our method obtain the prompted token embeddings $\mathbf{T}^{G}$ by extracting the structural information among tokens to prompt the pre-trained fundation models. Furthermore, the prompted token embeddings $\mathbf{T}^{G}$ are fed into a unified transformer-based encoder $\mathcal{F}_{uni}(\cdot)$ to obtain the representation of the subject, i.e., $\mathbf{z}=\mathcal{F}_{uni}(\mathbf{T}^{G}),\mathbf{z}\in\mathbb{R}^{d}$.
@@ IV-A1 Datasets
We conduct experiments on two public multimodal neurological disorder datasets, i.e., Alzheimer’s Disease Neuroimaging Initiative (ADNI) [^1] https://adni.loni.usc.edu/ (Jack Jr et al., 2008) and Autism brain imaging data exchange (ABIDE) [^2] http://fcon_1000.projects.nitrc.org/indi/abide/ (Di Martino et al., 2014). The subjects are categorized into four groups in ADNI: NC (normal control), EMCI (early mild cognitive impairment), LMCI (late mild cognitive impairment), and AD (Alzheimer’s disease). In total, we included 409 pairs of data (i.e., CN/LMCI/AD for ADNI-3CLS) and 770 pairs of data (i.e., CN/EMCI/LMCI/AD for ADNI-4CLS). The subjects are categorized into two groups in ABIDE: NC (normal control) and ASD (autism spectrum disorder) and we included 1029 pairs of data in ABIDE dataset.
We utilized the fslreorient2std, robustfov, FLIRT, and BET tools from the FSL software (Jenkinson et al., 2012) to preprocess data, the pipline are follows. We fistly employed the fslreorient2std tool to reorient the images, aligning them with the orientation of the standard template image. Next, we utilized the robustfov tool to crop the MRI images, effectively removing the neck and lower jaw regions. The FLIRT tool (Jenkinson et al., 2002) was used to register all MRI images to the Colin27 template (Holmes et al., 1998), correcting for global linear differences and resampling the images to a consistent resolution (1 × 1 × 1mm³) and size (191 × 217 × 191). Finally, we applied the BET tool to remove the skull and dura mater.
@@ IV-A2 Comparison methods
The comparison methods include one baseline transformer method (i.e., ViT (Dosovitskiy et al., 2020)), two transformer-based methods for neurological disorders diagnosis (i.e., BrainT (Kan et al., 2022) and GraphT (Cai et al., 2022)), two vanilla multi-modal large models (i.e., CLIP (Radford et al., 2021) and MetaT (Zhang et al., 2023a)), and two prompt learning methods (i.e., VPT (Jia et al., 2022) and MaPLe (Khattak et al., 2023)). We list the details of the comparison methods as follows:
  • ViT serves as the foundational transformer method. ViT applies the core ideas of transformers to image classification tasks. It treats images as sequences of patches and processes them using self-attention mechanisms, setting a precedent for subsequent transformer-based models in computer vision.
  • BrainT leverages transformer to efficiently learn connection strengths between brain regions and incorporating an orthonormal clustering readout operation to capture functional modules within the brain.
@@
  • CLIP is a multi-modal large model, which learning visual concepts from natural language supervision. It has been shown to be effective in a wide range of visual tasks by leveraging the power of language-image pre-training.
@@
  • VPT introduces an efficient prompt learning to full fine-tuning for large-scale transformer models in computer vision. This groundbreaking methodology not only streamlines the training process but also significantly optimizes the utilization of computational resources.
@@ IV-A3 Implementation details
All experiments are conducted on a server with 8 NVIDIA GeForce 3090 GPU (24.0GB caches). In our method, we utilize the text encoder and vision encoder from the pre-trained multimodal model CLIP as the initial weights for our text encoder and unified encoder. To achieve efficient fine-tuning, the parameters of these encoders do not need to be updated (i.e., frozen) during the fine-tuning process. Furthermore, this ensures that the embeddings obtained from the text encoder and unified encoder are aligned with each other. Additionally, all parameters are initialized by the Glorot initialization and optimized by the AdamW optimizer with initial learning rate $0.0001$, and decay it by 0.2 at epochs 30 and 60. MMGPL is trained with a fixed epoch 100. The patch size is set in the range of $\{16,24,32,64\}$ and the batch size is set in the range of $\{4,8,16\}$ according to memory. We obtained the author-verified codes for all comparison methods and tune the parameters for all comparison methods as suggested in their published works. Furthermore, for methods that are originally single-modal, we adopt the results from the best-performing modality as the final outcome. In the case of BrainT, we replace its tokenizer module to adapt to the modality data utilized in the experiment.
@@ IV-B1 Diagnose performance
Figure 4: Heat maps generated by MMGPL on different subjects in ADNI dataset, i.e., AD (ID: 018-S-0286), NC (ID: 007-S-1222), EMCI (ID: 002-S-0729), and LMCI (ID: 002-S-0729).
@@ IV-B2 Ablation studies
We further conduct the ablation studies with individual components and report their results in Fig. 2. Specifically, “B” denotes baseline method which apply the fine-turning strategy introduced in method MetaT (Zhang et al., 2023a). Based on this, we apply concept learning (i.e., Sec. III-C) to reduce the weights of irrelevant patches and denote as “B+C”. Similarly, we employ graph prompt learning (i.e., Sec. III-D) and denote as “B+G”. Note that, in order to ensure the proper running of the method, the semantic similarity is still need to calculated in method “B+G”. Moreover, we involve both concept learning and graph prompt learning which denotes as “B+C+G”.
Based on the experimental results, we could observe that each components has a contribution. In particular, compared to the baseline method “B”, the method “B+G” averagely improves by about 5%. Moreover, the method “B+C” markedly outperform that of the the baseline method “B”. Hence, the proposed concept learning and graph prompt learning modules have positive effect on the diagnosis of neurological disorder, which versified our claim that reduce the weight of irrelevant patches and incorporate a graph prompt making our method particularly suitable for neurological disorders diagnosis. Besides, we found that concept learning contributes more significantly to the overall performance compared to graph prompt. One possible reason is that without concept learning, the inclusion of many irrelevant tokens during graph prompt learning could have a negative impact. Considering the message propagation mechanism of GCN, these irrelevant tokens can potentially affect the graph prompt learning process and consequently affect the overall performance. Therefore, the two modules, concept learning and graph prompt learning, are closely intertwined and mutually dependent on each other.
@@ IV-B3 Analysis
Multi-modalities. We investigate the effectiveness of individual modality and multi-modalities by reporting the performances in Fig. 2. Concretely, our method considering multi-modalities (e.g., MRI and PET in ADNI) obtains substantial improvements across all metrics, compared to only considering individual modality. This suggests that the complementary information captured by different modalities can enhance the overall performance of the model.
@@
Interpretability. We further investigate the interpretability of our method. To achieve this, we scale the heat maps and superimpose them on the corresponding positions of the original images. The visualization of the heat map is presented in Fig. 4, with darker shades of red signifying higher weights and darker shades of blue signifying lower weights. Specifically, the x-axis represents different positions within the same plane, while the y-axis represents different planes. Firstly, it is evident that our method effectively eliminates the weights associated with almost all irrelevant patches. Secondly, Fig. 4 indicates that our method is capable of identifying the critical regions within the brain, such as hippocampal and parahippocampal. Morevoer, its interpretability has been validated by clinicians.
@@
Scalability. We further conduct comparison between MMGPL and related works in term of scalability. As shown in Table II. we evaluate three aspects of scalability, i.e., multiple modalities, unified model, and prompt. In particular, considering the inherent multimodality of medical data, we first evaluate the model’s ability to handle multiple modalities. Note that, $\checkmark$(vanilla) indicates can only supports two modalities (e.g., image and text) and is challenging to expand to supports more modalities. Secondly, we evaluate whether the method applies a unified model for encoding multimodal medical imaging, as employing a unified model facilitates efficient inference and deployment. Thirdly, we evaluate whether the method applies prompt for fine-turning, as prompt driven fine-turning is beneficial for enhancing the interactivity and effectiveness of the method. Comparing MMGPL with all comparison methods, we can observe that our method achieve excellent scalability and flexibility.
@@ V Conclusion
In this paper, we proposed a graph prompt learning fine-turning framework for neurological disorder diagnosis, by jointly considering the impact of irrelevant patches as well as the structural information among tokens in multimodal medical data. Specifically, we conduct concept learning, aiming to reduce weights of irrelevant tokens according to the semantic similarity between each token and disease-related concepts. Moreover, we conducted graph prompt learning with concept embeddings, aiming to bridge the gap between multimodal large models and neurological disease diagnosis. Experimental results demonstrated the effectiveness of our proposed method, compared to state-of-the-art methods on neurological disease diagnosis tasks.

### v2

@@ III-D2 Graph embedding
where $\tilde{\mathbf{T}}$ is the weighted embeddings of the tokens. As a result, our method obtains the prompted token embeddings $\mathbf{T}^{G}$ by extracting the structural information among tokens to prompt the pre-trained fundation models. Furthermore, the prompted token embeddings $\mathbf{T}^{G}$ are fed into a unified transformer-based encoder $\mathcal{F}_{uni}(\cdot)$ to obtain the representation of the subject, i.e., $\mathbf{z}=\mathcal{F}_{uni}(\mathbf{T}^{G}),\mathbf{z}\in\mathbb{R}^{d}$.
@@ IV-A1 Datasets
We conduct experiments on two public multimodal neurological disorder datasets, i.e., Alzheimer’s Disease Neuroimaging Initiative (ADNI) [^1] https://adni.loni.usc.edu/ (Jack Jr et al., 2008) and Autism brain imaging data exchange (ABIDE) [^2] http://fcon_1000.projects.nitrc.org/indi/abide/ (Di Martino et al., 2014). The subjects are categorized into four groups in ADNI: NC (normal control), EMCI (early mild cognitive impairment), LMCI (late mild cognitive impairment), and AD (Alzheimer’s disease). In total, we included 409 pairs of data (i.e., NC/LMCI/AD for ADNI-3CLS) and 770 pairs of data (i.e., NC/EMCI/LMCI/AD for ADNI-4CLS). The subjects are categorized into two groups in ABIDE: NC (normal control) and ASD (autism spectrum disorder) and we included 1029 pairs of data in ABIDE dataset.
We utilized the fslreorient2std, robustfov, FLIRT, and BET tools from the FSL software (Jenkinson et al., 2012) to preprocess data, the pipeline are follows. We first employed the fslreorient2std tool to reorient the images, aligning them with the orientation of the standard template image. Next, we utilized the robustfov tool to crop the MRI images, effectively removing the neck and lower jaw regions. The FLIRT tool (Jenkinson et al., 2002) was used to register all MRI images to the Colin27 template (Holmes et al., 1998), correcting for global linear differences and resampling the images to a consistent resolution (1 × 1 × 1mm³) and size (191 × 217 × 191). Finally, we applied the BET tool to remove the skull and dura mater.
@@ IV-A2 Comparison methods
The comparison methods include one baseline transformer method (i.e., ViT (Dosovitskiy et al., 2020)), two transformer-based methods for neurological disorders diagnosis (i.e., BrainT (Kan et al., 2022) and GraphT (Cai et al., 2022)), two vanilla multi-modal large models (i.e., CLIP (Radford et al., 2021) and MetaT (Zhang et al., 2023b)), and two prompt learning methods (i.e., VPT (Jia et al., 2022) and MaPLe (Khattak et al., 2023)). We list the details of the comparison methods as follows:
  • ViT is treated as a baseline method. ViT applies the core ideas of transformers to image classification tasks. It treats images as sequences of patches and processes them using self-attention mechanisms, setting a precedent for subsequent transformer-based models in computer vision.
  • BrainT leverages the transformer to efficiently learn connection strengths between brain regions and incorporate an orthonormal clustering readout operation to capture functional modules within the brain.
@@
  • CLIP is a multi-modal large model, which learns visual concepts from natural language supervision. It has been shown to be effective in a wide range of visual tasks by leveraging the power of language-image pre-training.
@@
  • VPT introduces efficient prompt learning to full fine-tuning for large-scale transformer models in computer vision. This groundbreaking methodology not only streamlines the training process but also significantly optimizes the utilization of computational resources.
@@ IV-A3 Implementation details
All experiments are conducted on a server with 8 NVIDIA GeForce 3090 GPU (24.0GB caches). In our method, we utilize the text encoder and vision encoder from the pre-trained multimodal model BiomedCLIP (Zhang et al., 2023a) as the initial weights for our text encoder and unified encoder. Moreover, BiomedCLIP was trained on 15 million pairs of medical image-text data. By leveraging the pre-trained BERT model from BiomedCLIP, we are able to effectively align concept texts with embeddings that possess semantic meaning in the image space. To achieve efficient fine-tuning, the parameters of these encoders do not need to be updated (i.e., frozen) during the fine-tuning process. Furthermore, this ensures that the embeddings obtained from the text encoder and unified encoder are aligned with each other. Additionally, all parameters are initialized by the Glorot initialization and optimized by the AdamW optimizer with initial learning rate $0.0001$, and decay it by 0.2 at epochs 30 and 60. MMGPL is trained with a fixed epoch 100. The patch size is set in the range of $\{16,24,32,64\}$ and the batch size is set in the range of $\{4,8,16\}$ according to memory. We obtained the author-verified codes for all comparison methods and tune the parameters for all comparison methods as suggested in their published works. Furthermore, for methods that are originally single-modal, we adopt the results from the best-performing modality as the final outcome. In the case of BrainT, we replace its tokenizer module to adapt to the modality data utilized in the experiment.
@@ IV-B1 Diagnose performance
Figure 4: Heat maps generated by MMGPL on different subjects in ADNI dataset.
@@ IV-B2 Ablation studies
We further conduct the ablation studies with individual components and report their results in Fig. 2. Specifically, “B” denotes the baseline method which applies the fine-turning strategy introduced in method MetaT (Zhang et al., 2023b). Based on this, we apply concept learning (i.e., Sec. III-C) to reduce the weights of irrelevant patches and denote as “B+W”. Similarly, we employ graph prompt learning (i.e., Sec. III-D) and denote as “B+G”. Note that, in order to ensure the proper running of the method, the semantic similarity is still needed to calculate in method “B+G”. Moreover, we involve both token weights and graph prompt learning which denotes as “B+W+G”.
Based on the experimental results, we could observe that each component has a contribution. In particular, compared to the baseline method “B”, the method “B+G” averagely improves by 5.3%. Moreover, the method “B+W” markedly outperforms that of the baseline method “B”. Hence, the proposed concept learning and graph prompt learning modules have a positive effect on the diagnosis of neurological disorders, which verified our claim that reduces the weight of irrelevant patches and incorporates a graph prompt making our method particularly suitable for neurological disorders diagnosis. Besides, we observed that concept learning with token weights contributes more significantly to the overall performance compared to graph prompts. One possible reason is that without token weights, the inclusion of many irrelevant tokens during graph prompt learning could have a negative impact. Considering the message propagation mechanism of GCN, these irrelevant tokens can potentially affect the graph prompt learning process and consequently affect the overall performance. Therefore, the two modules, concept learning with token weights and graph prompt learning, are closely intertwined and mutually dependent on each other.
@@ IV-B3 Analysis
Multi-modalities. We investigate the effectiveness of individual modalities and multi-modalities by reporting the performances in Fig. 2. Concretely, our method considering multi-modalities (e.g., MRI and PET in ADNI) obtains substantial improvements across all metrics, compared to only considering individual modalities. This suggests that the complementary information captured by different modalities can enhance the overall performance of the model.
@@
Interpretability.We further investigate the interpretability of our method. To achieve this, we scale the heat maps and superimpose them on the corresponding positions of the original images. signifying higher weights and darker shades of blue signifying lower weights. Specifically, the x-axis represents different positions within the same plane, while the y-axis represents different planes. Firstly, it is evident that our method effectively eliminates the weights associated with almost all irrelevant patches. Secondly, Fig. 4 indicates that our method is capable of identifying the critical regions within the brain, such as hippocampal and parahippocampal. Moreover, its interpretability has been validated by clinicians.
@@
Scalability. We further conduct a comparison between MMGPL and related works in terms of scalability. As shown in Table II. we evaluate three aspects of scalability, i.e., multiple modalities, unified model, and prompt. In particular, considering the inherent multimodality of medical data, we first evaluate the model’s ability to handle multiple modalities. Note that, $\checkmark$(vanilla) indicates can only support two modalities (e.g., image and text) and is challenging to expand to support more modalities. Secondly, we evaluate whether the method applies a unified model for encoding multimodal medical imaging, as employing a unified model facilitates efficient inference and deployment. Thirdly, we evaluate whether the method applies prompt for fine-turning, as prompt-driven fine-turning is beneficial for enhancing the interactivity and effectiveness of the method. Comparing MMGPL with all comparison methods, we can observe that our method achieves excellent scalability and flexibility.
(a) NC
(b) LMCI
(c) AD
Figure 5: The visualization of concept-similarity graph on the ADNI dataset. The horizontal and vertical axes represent concepts and tokens. Different colors represent concepts belonging to different categories. The red texts represent concepts related to NC, the green texts represent concepts related to LMCI, and the blue texts represent concepts related to AD.
Figure 6: The visualization of the quantified impact of different concepts on the ADNI dataset. The concepts are shown on the left side, while classes are shown on the right side. The width of the lines corresponds to the magnitude of the weights, and the values indicate the specific weight values.
Semantic similarity graph. To improve the interpretability of MMGPL, we visualize the semantic similarity graph of samples belonging to different classes in Fig. 5. Specifically, the semantic similarity graph is sparse. The reason could be that those concepts naturally have low similarity or limited overlap. Moreover, sparse semantic similarity graphs are easier to interpret as they highlight the most significant concepts. For example, in the middle sub-figure, the model infers the sample as LMCI category based on the fifth concept.
Concept weights. To quantify the weights of concepts, we conducted a statistical analysis on the number of times concepts were activated by samples belonging to different classes, and visualized the results using the Sankey diagram in Fig. 6. Firstly, the weights of concepts can accurately represent their corresponding categories. In the top side of Fig. 6, it can be observed that the top four concepts, which represent NC-related concepts, have the highest probability of being activated in NC samples. Moreover, concepts related to hippocampal atrophy tend to have higher weights and proportions in both LMCI and AD categories. This finding aligns with prior knowledge that hippocampal atrophy is associated with cognitive decline in patients with LMCI and AD. Secondly, There may be instances where samples from one category activate concepts belonging to other categories. For example, LMCI samples might activate concepts associated with the NC category. This could be due to individual variations, where pathological samples may not satisfy all the concepts. It is worth noting that NC samples rarely activate concepts related to AD, while AD samples may activate concepts associated with NC. This finding aligns with clinical expectations, as NC samples typically do not exhibit any AD-related symptoms. However, some AD samples may still satisfy some concepts associated with NC. Moreover, LMCI samples can originate from concepts related to both NC and AD. This finding aligns with clinical expectations, as LMCI represents an intermediate state between NC and AD.
@@ V Conclusion
In this paper, we proposed a graph prompt learning fine-turning framework for neurological disorder diagnosis, by jointly considering the impact of irrelevant patches as well as the structural information among tokens in multimodal medical data. Specifically, we conduct concept learning, aiming to reduce the weights of irrelevant tokens according to the semantic similarity between each token and disease-related concepts. Moreover, we conducted graph prompt learning with concept embeddings, aiming to bridge the gap between multimodal models and neurological disease diagnosis. Experimental results demonstrated the effectiveness of our proposed method, compared to state-of-the-art methods on neurological disease diagnosis tasks.